# Motion Generation from Fine-grained Textual Descriptions

### v1

@@ 3.3.   Human Evaluation of FineHumanML3D
We classify all text-motion pairs into three categories: zero, partial and perfect alignment. Further, the errors in partial alignment are classified into inversion (e.g., left leg in the text, right leg in the motion), mismatch (partial texts not matching motion clips), redundancy (stating more than the motion performs), and deficiency (lacking descriptions of motion clips). We also incorporate noticeable sub-errors of redundancy and deficiency. The beginning pose sub-error denotes the motion does not follow the beginning pose description, so for the ending pose sub-error. The insufficient repetition sub-error denotes the textual description decreases the repetition times of a single motion clip. Finally, counts of zero, partial and perfect alignment turn out to be 2 : 68 : 30. Concrete errors in partial alignment are shown in Table 2  [^5] Each text-motion pair might have several errors and sub-errors..
Despite the disagreement between fine-grained descriptions and golden motions, we are curious about to what extent a text2motion model trained on such data can correctly learn mappings from descriptions to motions, which leads to work below.
4.   Our Model
We design a new text2motion model, FineMotionDiffuse, which is composed of a fine-text encoder, a step-aware self-attention block, a coarse-text encoder, a fine-coarse cross-attention block and a diffusion block, to conduct training and inference on FineHumanML3D. An overview of our FineMotionDiffuse model is shown in Figure 3.
The fine-text encoder takes fine-grained motion descriptions with named step marks as input, and outputs step embeddings encoding full features of each step. In order to better capture the order among steps, we propose a new step-aware self-attention, which adds hard positional embeddings to the step embeddings and then passes them through one self-attention block. The coarse-text encoder digests the coarse-grained description and outputs the corresponding token embeddings. Outputs from the fine-text and coarse-text encoder are then put into the fine-coarse cross-attention block. For the diffusion block, We utilize the diffusion model architecture from Zhang et al. (2022) to fuse textual and motion features, and conduct training and inference.
@@ 4.   Our Model
Given step embeddings, our step-aware self-attention mechanism will add hard positional embeddings ($PE$) to them through sine and cosine functions from Vaswani et al. (2017), and feed these position-encoded step embeddings into a self-attention Transformer block ($SelfAtt$) to get fine-grained textual features ($f\_feas$). This mechanism helps better capture the temporal relationships among steps, and improve the quality of fine-text features through interactions among step embeddings.
@@
5.   Experiments
We design our experiments to answer the following questions: 1) What are the effects of different blocks and mechanisms in FineMotionDiffuse? 2) What advantages does FineMotionDiffuse own over MotionDiffuse? 3) Can FineMotionDiffuse deal with composite motions?
5.1.   Settings
@@
We use the 85,646 fine-grained motion descriptions along with the 19,228 motions from our FineHumanML3D dataset. For coarse-grained motion descriptions, we use the counterparts of our fine-grained descriptions in the original HumanML3D dataset, counting to 85,646. The training dataset consists of 23,966 motions with 70,222 descriptions, while the testing dataset consists of 5,262 motions with 15,424 descriptions.
Evaluation Metrics.
Following Guo et al. (2022b), we pretrain a text-motion contrastive model on the training set of FineHumanML3D for evaluation purpose. With the text encoder and motion encoder from this model, we apply three evaluation metrics. 1) Frechet Inception Distance (FID): The generated motion and its ground-truth motion are passed through the motion encoder, and FID is the distance between the two encoded motion sequences. 2) R-precision: For one text and its generated motion sequence, 31 mismatched texts are sampled from the testing set. The motion sequence is encoded by the motion encoder, and 32 texts are encoded by the text encoder. The top-k (k=1,2) accuracy is calculated by checking whether the true text falls within the top-k nearest texts from the motion. 3) Diversity: All generated motion sequences in the testing set are randomly split into pairs, and the average joint difference is calculated as the diversity metric.
Baseline.
We propose MotionDiffuse_AddFC as our baseline model. It first respectively encodes the coarse-grained description and each step in the fine-grained description with CLIP, which results in $1+n$ tensors of the size (context_length, token_embedding_dim). Then we directly add the tensors at each token position, resulting in one fine-coarse feature tensor.
5.2.   Main Results
The implementation details of main experiments are in Appendix B. As shown in Table 3, while the two models are comparable in R-precision, FineMotionDiffuse performs significantly better in FID. This indicates the hierarchical attention architecture helps FineMotionDiffuse produce more accurate motion sequences, aligning more precisely with the ground-truth motions. Additionally, FineMotionDiffuse’s improvement in Diversity shows its stronger ability to recognize details from fine-grained descriptions.
Methods FIDnormal-$\downarrow$ R-precisionnormal-$\uparrow$ Diversitynormal-$\rightarrow$
@@ 5.2.   Main Results
5.3.   Ablation Study
For ablation study, we investigate the following variants of our method: 1) unfreezing the CLIP parameters, 2) choices for step representations, 3) the effect of fine-coarse cross-attention, and 4) truncating input step texts.
Due to high GPU consumption when training and evaluating on the whole dataset, we take 2,046 motions for training from the whole training set, and 878 motions for evaluation from the whole testing set. Table 4 reports quantitative evaluation results of these ablated variants, where FID is taken for comparison.
Model Ablation.
After we unfreeze parameters in CLIP ($+$ CLIP_unfrozen), FID increases by 0.13. We guess that freezing these parameters might preserve the generalization ability of the model.
Next, we take the embedding of [E] token of each step ($\rightarrow$[E]), instead of averaging along all token embeddings. FID increases by 0.5, showing the advantage of acquiring a step embedding by taking average. Further, we remove the cross-attention with coarse-grained descriptions (-- cross-attention), and find that FID increases by a large margin of 1.2. This proves that the cross-attention between fine-grained and coarse-grained descriptions greatly helps the model generate more precise motions.
Input Ablation.
@@ 5.3.   Ablation Study
5.4.   Qualitative Analysis
We aim to compare motions generated by FineMotionDiffuse and MotionDiffuse, by investigating: 1) How well they conform to the input descriptions; 2) How well motion details are performed (e.g., whether feet slide on the ground, whether body parts shake frequently, etc.). Therefore, we collect 30 coarse-grained motion descriptions (9 for seen basic motions and 21 for unseen composite motions), and expand them into fine-grained ones with GPT-3.5-turbo, using our best prompt.
For each motion type, we generate three motion sequences with three methods: 1) Feeding the fine-grained description along with the coarse-grained one into FineMotionDiffuse; 2) Feeding the coarse-grained description into MotionDiffuse (MotionDiffuse_coarse); 3) Feeding the fine-grained description (named step marks removed) into MotionDiffuse (MotionDiffuse_detailed).
We conduct human evaluation on all generated sequences. Each sequence is scored from 1 to 5, with 0.5 as a unit. We respectively calculate the average scores of all motion sequences, basic motion sequences and composite motion sequences for each method. Additionally, we also report the score drops from basic to composite motions ($\Delta$). Results are shown in Table 5.
Meanwhile, FineMotionDiffuse performs better than MotionDiffuse_coarse, which proves motion details are more effectively learned during its training. Furthermore, FineMotionDiffuse has the smallest score drop from basic to composite motions, showing stronger generalization ability to deal with unseen composite motion descriptions.
Methods allnormal-$\uparrow$ basicnormal-$\uparrow$ compositenormal-$\uparrow$ $\Delta\downarrow$
@@ 5.4.   Qualitative Analysis
Table 5: Human evaluation on 30 new motions.
5.5.   Case Study
Here we investigate model performance on composite motions from two dimensions, spatial and chronological compositionality. Spatial compositionality denotes the model’s ability to simultaneously perform two basic motions, and chronological compositionality indicates performing two basic motions in a sequential manner. We choose two groups of cases for qualitative analysis. Full details of the input descriptions are in Appendix C.
@@ 5.5.   Case Study
While FineMotionDiffuse has respectively learned how to slightly squat (Figure 4(a)) and raise both arms above the head (Figure 4(b)), it is able to compose the corresponding composite motion that simultaneously performs the two basic motions (Figure 4(c)). On the contrast, MotionDiffuse fails to handle motions of arms when dealing with fine-grained texts (Figure 4(d)). Taking the coarse-grained texts as input, MotionDiffuse does not show the process of squatting down in the beginning (Figure 4(e)).
@@
Figure 4: Example motions with spatial compositionality. $\color[rgb]{1,0,0}C$ denotes inputting coarse-grained texts, $\color[rgb]{1,0,0}F$ denotes inputting fine-grained texts, and $\color[rgb]{1,0,0}C+F$ denotes inputting both texts. Due to space limit, we only display coarse-grained descriptions here. Fine-grained descriptions are shown in Appendix C.
@@
While FineMotionDiffuse has respectively learned how to walk (Figure 5(a)) and kick with one leg (Figure 5(b)), it successfully composes the corresponding composite motion that concatenate the two basic motions in time order (Figure 5(c)). On the contrast, MotionDiffuse fails to kick when dealing with fine-grained texts (Figure 5(d)). Taking the coarse-grained texts as input, MotionDiffuse fails to make the agent walk(Figure 5(e)).
@@
Figure 5: Examples with chronological compositionality. Fine-grained descriptions are in Appendix C.
@@
In the two groups of cases, FineMotionDiffuse successfully shows the ability of spatial and chronological compositionality, while MotionDiffuse fails. On the one hand, MotionDiffuse trained on large amounts of coarse-grained texts does not learn mappings from fine-grained motion-related words to motion primitives, which explains its failure in spatial and chronological composition when dealing with fine-grained texts. On the other hand, when the inputs are coarse-grained composite descriptions, MotionDiffuse cannot show the whole motions. It might only learn how to map a seen coarse text into a single motion, but fails in acquiring the ability of generating an unseen composite motion from combination of seen coarse texts.
6.   Conclusion
In this paper, we design dedicate prompts to drive GPT-3.5-turbo to build a large-scale language-motion dataset with fine-grained textual descriptions, paving the road to investigate fine-grained text2motion. We find a step-by-step instruction with a pseudo-code compulsory check can greatly improve the generated descriptions. We propose FineMotionDiffuse to make best of both fine- and coarse-grained descriptions. Both quantitative and qualitative evaluations show that our method can better learn the implicit mappings from fine-grained descriptions and exhibit stronger generalization ability compared to previous state-of-the-art models, especially dealing with complex spatial and chronological composition motions.
@@ 6.   Conclusion
Dataset. The fine-grained descriptions in FineHumanML3D come from GPT-turbo-3.5 pretrained on massive amounts of texts and codes. Though they learned lots of knowledge about human bodies and physical motions, the direct expansion from coarse-grained motion descriptions still unavoiably deviates from the golden motion sequences, which, to some degree, leads to misbehaviour of the newly trained FineMotionDiffuse. We expect future work to automatically check and correct the generated descriptions, so that they can conform to the ground-truth motions. Also, large pretrained multimodal models like GPT-4 (OpenAI, 2023) can be used for dataset construction. Even more, the golden motions may be converted to some kinds of LLM-friendly inputs, so that LLM directly receives exact motion information, which helps generate more high-quality fine-grained descriptions.
Model. The current diffusion-based model architecture is quite slow in both training and inference, which hinders its application in real-world scenarios. Therefore, several acceleration methods (Yang et al., 2022; Ulhaq et al., 2022) can be adopted to alleviate this problem. Moreover, as some VAE-based models (Zhang et al., 2023) are reported to acquire comparable performances to diffusion-based ones, our method also shows potential in application to such architectures.
@@
The original dataset we use, HumanML3D, is not published directly, due to the distribution policy of one of its origins, AMASS dataset. Therefore, we follow the official instructions from the builders of HumanML3D (https://github.com/EricGuo5513/HumanML3D) to reproduce this dataset. Through our work, we only call OpenAI’s public API for GPT-3.5-turbo-0301. All experiments along with the dataset construction cost less than \$350. The text-tokenizer and POS tagger we use come from an open-source python package Spacy3.5. We recruit our colleagues to conduct human scoring and fairly pay them with more than \$10 per hour.
7.   Bibliographical References
[…]
@@ Appendix A Our Prompts and Evaluation
P1
[system] Please provide a detailed expansion of the following action description. The new description should be in chronological order. It should specify spatial position changes (including angle changes) of relevant body parts and muscle tension.
@@
P2
[system] Please provide a detailed expansion of one coarse-grained motion description. The new description should be in chronological order. It should specify spatial position changes (including angle changes) of relevant body parts.
@@
P3
[system] Please provide a detailed expansion of the following action description. The new description should be in chronological order. It should specify spatial position changes (including angle changes) of relevant body parts. Here is one example. You should first generate thinking, then the paraphrased description.
@@
P4
[system] Please provide a detailed expansion of one coarse-grained motion description. The new description should be in chronological order and step by step. It should specify spatial position changes (including angle changes) of relevant body parts.
@@
P5
[system] Please provide a detailed expansion of one coarse-grained motion description. The new description should be in chronological order and step by step. It should specify spatial position changes (including angle changes) of relevant body parts.
@@
P6
[system] In the first paragraph, please provide a detailed expansion of one coarse-grained motion description. The new description should be in chronological order and step by step. It should specify spatial position changes (including angle changes) of relevant body parts. It should not specify any information related to muscles. In the second paragraph, please convert the description in the first paragraph into a pseudo-code format.
@@
P7
[system] In the first paragraph, please provide a detailed expansion of one coarse-grained motion description. The new description should be in chronological order and step by step. It should specify spatial position changes (including angle changes) of relevant body parts. It should not specify any information related to muscles. In the second paragraph, please convert the description in the first paragraph into a pseudo-code format.
@@
P8
[system] In the first paragraph, please provide a detailed expansion of one coarse-grained motion description. The new description should be in chronological order and step by step. It should specify spatial position changes (including angle changes) of relevant body parts. It should not specify any information related to muscles. In the second paragraph, please convert the description in the first paragraph into a pseudo-code format.
@@
P1
Too many details other than body movements;
@@
P3
Stronger time order;
May sometimes be too detailed (e.g., finger tips);
@@
P4
Much more concise, barely with unnecessary statements;
@@
Table 7: The pilot prompts we test for text expansion from coarse-grained motion descriptions to fine-grained ones. We test each of them on 10 example coarse-grained descriptions, evaluate fine-grained descriptions returned by GPT-3.5-turbo, and give a qualitative analysis of pros and cons. The pros of each prompt come from comparisons with all prompts above it.
@@
Appendix B Implementation Details
Appendix C Input Texts for Case Studies

### v2

@@ 3.3 Human Evaluation of FineHumanML3D
While only 30 out of 100 motions are perfectly aligned with the fine-grained descriptions, perfect alignment is an exceptionally high bar. In the majority of cases, partial alignment indeed captures the correct time order and relationships among core motions. Issues like redundancy or deficiency are often trivial in nature. Substantive errors rarely occur outside of very complicated motions, e.g., juggling while doing the moonwalk. Therefore, despite the partial disagreement between fine-grained descriptions and golden motions, we aim to verify the usefulness and effectiveness of our data generation method through the performance of the downstream text2motion task.
4 Our Model
We design a new text2motion model, FineMotionDiffuse, which is composed of a fine-text encoder, a step-aware self-attention block, a coarse-text encoder, a fine-coarse cross-attention block, and a diffusion block, to conduct training and inference on FineHumanML3D. An overview of our FineMotionDiffuse model is shown in Figure 3.
The fine-text encoder takes fine-grained motion descriptions with named step marks as input, and outputs step embeddings encoding the full features of each step. In order to better capture the order among steps, we propose a new step-aware self-attention, which adds hard positional embeddings to the step embeddings and then passes them through one self-attention block. The coarse-text encoder digests the coarse-grained description and outputs the corresponding token embeddings. Outputs from the fine-text and coarse-text encoder are then put into the fine-coarse cross-attention block. For the diffusion block, We utilize the diffusion model architecture from Zhang et al. (2022) to fuse textual and motion features, and conduct training and inference.
@@ 4 Our Model
Given step embeddings, our step-aware self-attention mechanism adds hard positional embeddings ($PE$) to them through sine and cosine functions from Vaswani et al. (2017), and feeds these position-encoded step embeddings into a self-attention Transformer block ($SelfAtt$) to get fine-grained textual features ($f\_feas$). This mechanism helps better capture the temporal relationships among steps, and improve the quality of fine-text features through interactions among step embeddings.
@@
5 Experiments
We design our experiments to answer the following questions: 1) What are the benefits of different blocks and mechanisms in FineMotionDiffuse? 2) What advantages does FineMotionDiffuse (trained on FineHumanML3D) own over MotionDiffuse (trained on HumanML3D)? 3) Can FineMotionDiffuse deal with spatially or chronologically composite motions?
5.1 Settings
@@
We use the 85,646 fine-grained motion descriptions along with the 29,228 motions from our FineHumanML3D dataset. For coarse-grained motion descriptions, we use the counterparts of our fine-grained descriptions in the original HumanML3D dataset, counting to 85,646. The training dataset consists of 23,966 motions with 70,222 descriptions, while the testing dataset consists of 5,262 motions with 15,424 descriptions.
[…]
@@
Evaluation Metrics.
Following Guo et al. (2022b), we pre-train a text-motion contrastive model on the training set of FineHumanML3D for evaluation purposes. With the text encoder and motion encoder from this model, we apply three evaluation metrics. 1) Frechet Inception Distance (FID): The generated motion and its ground-truth motion are passed through the motion encoder, and FID is the distance between the two encoded motion sequences. 2) R-precision: For one text and its generated motion sequence, 31 mismatched texts are sampled from the testing set. The motion sequence is encoded by the motion encoder, and 32 texts are encoded by the text encoder. The top-k (k=1,2) accuracy is calculated by checking whether the true text falls into the top-k nearest texts from the motion. 3) Diversity: All generated motion sequences in the testing set are randomly split into pairs, and the average joint difference is calculated as the diversity metric.
Baseline.
We build on MotionDiffuse (Zhang et al., 2022) to propose MotionDiffuse_AddFC as our baseline model, following an intuitive idea to fuse fine-grained textual features by simply adding tensors of each step embedding. It first respectively encodes the coarse-grained description and each step in the fine-grained description with CLIP, which results in $1+n$ tensors of the size (context_length, token_embedding_dim). Then we directly add the tensors at each token position, resulting in one fine-coarse feature tensor.
Implementation Details.
5.2 Main Results
As shown in Table 3, while the two models are comparable in R-precision, FineMotionDiffuse significantly improves FID by 0.38. This indicates the hierarchical attention architecture helps FineMotionDiffuse produce more accurate motion sequences, aligning more precisely with the ground-truth motions. Additionally, FineMotionDiffuse’s improvement in Diversity shows its stronger ability to recognize details from fine-grained descriptions.
5.3 Ablation Study
For the ablation study, we investigate the following variants of our method: 1) Unfreezing the CLIP parameters; 2) Different choices of step representations; 3) The effect of fine-coarse cross-attention; 4) Truncating input step texts.
@@ 5.3 Ablation Study
Due to high GPU consumption when training and evaluating the whole dataset, we take 2,046 motions for training from the whole training set, and 878 motions for evaluation from the whole testing set. Table 4 reports quantitative evaluation results of these ablated variants, where FID is taken for comparison.
Model Ablation.
After we unfreeze parameters in CLIP ($+$ CLIP_unfrozen), FID increases by 0.13. We guess that freezing these parameters preserves the generalization ability of the model.
Next, we take the embedding of [E] token of each step ($\rightarrow$[E]), instead of averaging along all token embeddings. FID increases by 0.5, showing the advantage of acquiring a step embedding by taking the average. Further, we remove the cross-attention with coarse-grained descriptions (-- cross-attention), and find that FID increases by a large margin of 1.2. This proves that the cross-attention between fine-grained and coarse-grained descriptions greatly helps the model generate more precise motions.
Input Ablation.
For the second group, we directly use our originally trained FineMotionDiffuse, but make some modifications when testing: 1) We remove the first and last steps (referred to as delFirstLastinput), and it is quite surprising to find that FID improves by 0.13; 2) We keep the first and last steps only when testing (referred to as delInnerinput), and find that it suffers from severe performance drops. Careful investigation into the fine-grained descriptions reveals that the first and last steps usually describe beginning and ending poses, which are mostly motionless states compared with actions in those inner steps, and do not vary much across different motions. Therefore, we think that FineMotionDiffuse trained on FineHumanML3D is more sensitive to rich actions in the inner steps, and its generalization ability tends to be disturbed by the relatively uninformative first and last steps.
5.4 Qualitative Analysis
We aim to compare motions generated by FineMotionDiffuse and MotionDiffuse, by investigating: 1) How well they conform to the input descriptions; 2) How well motion details are performed, e.g., whether feet slide on the ground, whether body parts shake frequently, etc. Therefore, we collect 30 coarse-grained motion descriptions (9 for seen basic motions and 21 for unseen composite motions), and expand them into fine-grained ones with the LLM, using our prompt template.
For each motion type, we generate three motion sequences with three methods: 1) Feeding the fine-grained description along with the coarse-grained one into FineMotionDiffuse; 2) Feeding the coarse-grained description into MotionDiffuse (MotionDiffuse_coarse); 3) Feeding the fine-grained description (named step marks removed) into MotionDiffuse (MotionDiffuse_detailed), with the texts being truncated due to the 77-token input length limit of CLIP.
We conduct human evaluation on all generated sequences. Each sequence is scored from 1 to 5, with 0.5 as a unit. We respectively calculate the average scores of all motion sequences, basic motion sequences, and composite motion sequences for each method. Additionally, we also report the score drops from basic to composite motions ($\Delta$). Results are shown in Table 5.
Methods $\uparrow$ $\uparrow$ $\uparrow$ $\Delta\downarrow$
@@ 5.4 Qualitative Analysis
Table 5: Human evaluation on 30 motions.
Meanwhile, FineMotionDiffuse performs better than MotionDiffuse_coarse, which proves motion details are more effectively learned during its training. Furthermore, FineMotionDiffuse has the smallest score drop from basic to composite motions, showing a stronger generalization ability to deal with unseen composite motion descriptions.
5.5 Case Study
Here we investigate model performance on composite motions from two dimensions, spatial and chronological compositionality. Spatial compositionality denotes the model’s ability to simultaneously perform two basic motions, and chronological compositionality indicates performing two basic motions in a sequential manner. We choose two groups of cases for qualitative analysis. Full details of the input descriptions are in Appendix B.
@@ 5.5 Case Study
While FineMotionDiffuse respectively learns how to slightly squat (Figure 4(a)) and raise both arms above the head (Figure 4(b)), it is able to compose the corresponding composite motion that simultaneously performs the two basic motions (Figure 4(c)). In contrast, MotionDiffuse fails to handle the motions of arms when dealing with fine-grained texts (Figure 4(d)). Taking the coarse-grained texts as input, MotionDiffuse does not show the process of squatting down in the beginning (Figure 4(e)).
@@
Figure 4: Example motions with spatial compositionality. $\color[rgb]{1,0,0}C$ denotes inputting coarse-grained texts, $\color[rgb]{1,0,0}F$ denotes inputting fine-grained texts, and $\color[rgb]{1,0,0}C+F$ denotes inputting both texts. Due to space limits, we only display coarse-grained descriptions here. Fine-grained descriptions are shown in Appendix B.
@@
While FineMotionDiffuse respectively learns how to walk (Figure 5(a)) and kick with one leg (Figure 5(b)), it successfully composes the corresponding composite motion that concatenates the two basic motions in time order (Figure 5(c)). In contrast, MotionDiffuse fails to kick when dealing with fine-grained texts (Figure 5(d)). Taking the coarse-grained texts as input, MotionDiffuse fails to make the agent walk (Figure 5(e)).
@@
Figure 5: Examples with chronological compositionality. Fine-grained descriptions are in Appendix B.
@@
In the two groups of cases, FineMotionDiffuse successfully shows the ability of handling spatial and chronological compositionalities, while MotionDiffuse fails. On the one hand, MotionDiffuse trained on large amounts of coarse-grained texts does not learn mappings from fine-grained motion-related descriptions to motion primitives, which explains its failure in spatial and chronological composition when dealing with fine-grained texts. On the other hand, when inputs are coarse-grained composite descriptions, MotionDiffuse cannot show the whole motions. It might only learn how to map a seen coarse text into a single motion, but fails in acquiring the ability to generate an unseen composite motion from the combination of seen coarse texts.
6 Conclusion
In this paper, we drive the LLM to build FineHumanML3D, the first large-scale language-motion dataset specializing in fine-grained textual descriptions, paving the road to investigate fine-grained text2motion. We find that a step-by-step instruction with a pseudo-code compulsory check can greatly improve the quality of generated descriptions. We propose FineMotionDiffuse to make the best of both fine- and coarse-grained descriptions. Both quantitative and qualitative evaluations show that FineMotionDiffuse trained on FineHumanML3D can better learn the mappings from fine-grained descriptions and exhibit stronger generalizability compared to previous state-of-the-art models, especially when dealing with spatially and chronologically composite motions. In the future, we hope to build on such methods to scale towards increasingly complex human behaviors.
@@ 6 Conclusion
Dataset. The fine-grained descriptions in FineHumanML3D come from an LLM pretrained on extensive texts and codes. Although the LLM learns lots of knowledge about human bodies and physical motions, the direct expansion from coarse-grained motion descriptions unavoidably deviates from the golden motion sequences, which, to some degree, leads to the misbehavior of newly trained FineMotionDiffuse. We expect future work to automatically check and correct the generated descriptions, so that they conform to the ground-truth motions. Also, large pre-trained multimodal models like GPT-4 (OpenAI, 2023) can be used for dataset construction. Even more, the golden motions may be converted to some kinds of LLM-friendly inputs, so that LLM directly receives exact motion information, which helps generate more high-quality fine-grained descriptions.
Model. The current diffusion-based model architecture is very slow in both training and inference, which hinders its application in real-world scenarios. Therefore, several acceleration methods (Yang et al., 2022; Ulhaq et al., 2022) can be adopted to alleviate this problem. Moreover, as some VAE-based models (Zhang et al., 2023) are reported to acquire comparable performance to diffusion-based ones, our method also shows potential in application to such architectures.
@@
The original dataset we use, HumanML3D, is not published directly, due to the distribution policy of one of its origins, AMASS dataset. Therefore, we follow the official instructions from the builders of HumanML3D (https://github.com/EricGuo5513/HumanML3D) to reproduce this dataset. Throughout the work, we only call OpenAI’s public API for GPT-3.5-turbo-0301. All experiments along with the dataset construction cost less than \$350. The text tokenizer and POS tagger we use come from an open-source Python package spaCy 3.5 (https://spacy.io/usage/v3-5). We recruit our colleagues to conduct human scoring and fairly pay them with more than \$10 per hour.
Acknowledgement
This work is supported by NSFC (62161160339). We would like to thank the anonymous reviewers for their helpful comments and suggestions. We would also like to thank Jason Narowdowsky for helping with the rebuttal and proofreading, and Yusuke Miyao for helping with the rebuttal.
@@ Appendix A Our Prompts and Evaluation
P1 [system] Please provide a detailed expansion of the following action description. The new description should be in chronological order. It should specify spatial position changes (including angle changes) of relevant body parts and muscle tension.
@@
P2 [system] Please provide a detailed expansion of one coarse-grained motion description. The new description should be in chronological order. It should specify spatial position changes (including angle changes) of relevant body parts.
@@
P3 [system] Please provide a detailed expansion of the following action description. The new description should be in chronological order. It should specify spatial position changes (including angle changes) of relevant body parts. Here is one example. You should first generate thinking, then the paraphrased description.
@@
P4 [system] Please provide a detailed expansion of one coarse-grained motion description. The new description should be in chronological order and step by step. It should specify spatial position changes (including angle changes) of relevant body parts.
@@
P5 [system] Please provide a detailed expansion of one coarse-grained motion description. The new description should be in chronological order and step by step. It should specify spatial position changes (including angle changes) of relevant body parts.
@@
P6 [system] In the first paragraph, please provide a detailed expansion of one coarse-grained motion description. The new description should be in chronological order and step by step. It should specify spatial position changes (including angle changes) of relevant body parts. It should not specify any information related to muscles. In the second paragraph, please convert the description in the first paragraph into a pseudo-code format.
@@
P7 [system] In the first paragraph, please provide a detailed expansion of one coarse-grained motion description. The new description should be in chronological order and step by step. It should specify spatial position changes (including angle changes) of relevant body parts. It should not specify any information related to muscles. In the second paragraph, please convert the description in the first paragraph into a pseudo-code format.
@@
P8 [system] In the first paragraph, please provide a detailed expansion of one coarse-grained motion description. The new description should be in chronological order and step by step. It should specify spatial position changes (including angle changes) of relevant body parts. It should not specify any information related to muscles. In the second paragraph, please convert the description in the first paragraph into a pseudo-code format.
@@
P1 - Too many details other than body movements;
@@
P3 Stronger time order; May sometimes be too detailed (e.g., finger tips);
@@
P4 Much more concise, barely with unnecessary statements;
@@
Table 7: The pilot prompts we test for text expansion from coarse-grained motion descriptions to fine-grained ones. We test each of them on 10 example coarse-grained descriptions, evaluate fine-grained descriptions returned by GPT-3.5-turbo-0301, and give a qualitative analysis of pros and cons. The pros of each prompt come from comparisons with all prompts above it.
@@
Appendix B Input Texts for Case Study